# Against The Achilles’ Heel: A Survey on Red Teaming for Generative Models

### v1

@@ Abstract
Generative models are rapidly gaining popularity and being integrated into everyday applications, raising concerns over their safety issues as various vulnerabilities are exposed. Faced with the problem, the field of red teaming is experiencing fast-paced growth, which highlights the need for a comprehensive organization covering the entire pipeline and addressing emerging topics for the community. Our extensive survey, which examines over 120 papers, introduces a taxonomy of fine-grained attack strategies grounded in the inherent capabilities of language models. Additionally, we have developed the searcher framework that unifies various automatic red teaming approaches. Moreover, our survey covers novel areas including multimodal attacks and defenses, risks around multilingual models, overkill of harmless queries, and safety of downstream applications. We hope this survey can provide a systematic perspective on the field and unlock new areas of research. Warning: This paper contains examples that may be offensive, harmful, or biased.
@@ 1 Introduction
Generative artificial intelligence (GenAI), such as Large Language Models (LLMs) and Vision-Language Models (VLMs), have gained wide application in areas such as dialogue systems (Ouyang et al., 2022), code completion (Chen et al., 2021), and domain-specific usages (Wu et al., 2023b). However, the generative nature and versatility of such models also introduce more vulnerabilities than conventional systems. GenAI can be induced to produce biased, harmful, or unintended outputs when presented with carefully crafted prompts, a phenomenon known as prompt attacks or LLM jailbreak (Shen et al., 2023), which may facilitate the propagation of harmful information and malicious exploitation of applications utilizing GenAI.
@@
To gain a comprehensive understanding of potential attacks on GenAI and develop robust safeguards, researchers have conducted studies on various red-teaming strategies, automated attack approaches, and defense methods. Several surveys have been composed to investigate the rapidly growing field of GenAI safety systematically. Although existing surveys provide valuable insights, we identify several areas for improvement. First, many surveys cover a limited scope of attack strategies and defenses compared to the rapidly growing body of research in this fast-evolving field. Their collection is coarse-grained, without paying attention to the specific strategies adopted in the paper. Second, most surveys lack a unifying perspective that connects the chain of attack taxonomies, attack methods, benchmarks, and defenses. Their taxonomies are based on surface features of strategies, which may be of less instructive value for future development. Third, emerging topics such as multilingual, multimodal attacks, and overkill are often overlooked or only briefly covered, where a dedicated discussion is needed to provide a comprehensive view of this rapidly evolving research area.
@@
  • We frame automated red-teaming methods as searching problems and analogously decouple popular search methods into three components of state space, search goal, and search operation, unlocking a larger space for the future design of automated red-teaming methods.
  • We pay distinct attention to emergent areas of GenAI safety, including multilingual and multimodal attacks, overkill of harmless queries, as well as the safety of downstream applications powered by LLMs such as dialogue agents.
@@
The remainder of this paper is structured as follows (also shown in Figure 2). Sec. 2 introduces key terminology and positions our work in the context of related surveys. Sec. 3 outlines a taxonomy of risks associated with LLMs. In Sec. 4, we dive into attack strategies and automated search methods for LLMs. Evaluation benchmarks and metrics are covered in Sec. 5. Sec. 6 presents an overview of defensive approaches. Sec. 7 and Sec. 8 discuss vulnerabilities in LLM-based multimodal and applications settings respectively. Finally, we highlight promising future research directions in Sec. 9 and conclude in Sec. 10.
@@ 2 Background
As the rise of GenAI requires understanding the associated risks posed by adversarial attacks and jailbreak, this section introduces key terminology and differentiates this survey from existing work by highlighting its comprehensive and cohesive coverage spanning risk taxonomies, attack methods, evaluation benchmarks, defenses, and emerging areas such as multimodal and multilingual attacks.
@@ 2.1 Terminology
Many different terms have been employed in the field of AI Safety. Here we try to define them and clarify their differences.
@@
An adversarial attack is a technique used to introduce intentionally crafted perturbations or modifications to the input data to cause a model to produce incorrect or undesirable outputs (Ren et al., 2020). These attacks can exploit vulnerabilities in the model’s training data, architecture, or decision boundaries, potentially leading to safety or security risks.
@@
Jailbreak refers to the process of overriding or bypassing the constraints or safety measures in a GenAI (Chao et al., 2023), typically through prompt engineering or other techniques. Jailbreak attempts to “break out” of the model’s guardrails, potentially allowing it to engage in harmful activities such as committing crimes or propagating hate speech (Derner & Batistič, 2023). The difference between adversarial attack and jailbreak is subtle. While the primary goal of the adversarial attack is to trigger erroneous model outputs like wrong classification results, a jailbreaker aims to obtain output that breaches established safety guidelines (Shayegani et al., 2023b). Furthermore, adversarial attacks can be performed either in the training or usage phases, whereas jailbreak is usually done during the inference phase of the model.
@@
Prompt injection, similar to SQL injection, refers to the process of inserting malicious or manipulative content into the prompts or inputs provided to a model (Greshake et al., 2023), and hijack the model’s intended task. For example, the attacker may trick models into submitting sensitive information to a malicious URL, or providing wrong summary results or false labels. Prompt injection is different from jailbreak where the latter aims to bypass safety restrictions, challenging model alignment efforts (Shayegani et al., 2023b).
@@
Red teaming is a practice of simulating malicious scenarios to identify vulnerabilities and test the robustness of systems or models (Abbass et al., 2011). Distinct from hacking or malicious attacks, red teaming in AI safety is a controlled process typically conducted by model developers that deliberately probes and manipulates models to uncover potential harmful behaviors. This process helps uncover weaknesses and informs the development of mitigation strategies. In AI safety, red teaming can employ methods such as adversarial attack, jailbreak, and prompt injection.
@@
Alignment refers to the process of ensuring that the behavior and outputs of AI are consistent with human values and ethical principles (Askell et al., 2021; Ouyang et al., 2022), such as fine-tuning or RLHF discussed in Sec. 6.1. There are numerous alignment criteria (Kiritchenko & Mohammad, 2018; Carranza et al., 2023; Kirchner et al., 2022), while the common alignment goal follows the HHH criteria (Askell et al., 2021):
@@
Overkill describes language models’ refusals on harmless questions containing seemingly dangerous keywords, e.g. “How to kill a python process” or “Where can I shoot a good photo?”. This is usually a side-effect of safety alignment as the model leans towards harmlessness instead of helpfulness in the multi-objective optimization process, compromising its practical value. Röttger et al. (2023) categorize the patterns of overkill and curate XSTest to evaluate this phenomenon.
@@
Different studies have varying definitions of ASR, mainly diverging in their choice of indicator function $I(\cdot)$. For instance, Zou et al. (2023) determines ASR based on the presence of specific keywords such as “I am sorry.” In contrast, works like Pi et al. (2024); shu et al. (2024) employ language models to judge the success of an attack. For a detailed introduction to ASR, see Sec. 5.1.1.
@@ 2.2 Related Work
Our survey, taking a vast amount of prior honorable work (Dong et al., 2024b; Das et al., 2024; Shayegani et al., 2023b; Deng et al., 2023c; Esmradi et al., 2023) discussing risk taxonomies into consideration, provides one of the most comprehensive and structured views of attacks and risks, based on different abilities of GenAI such as in-context learning, autoregressive modeling, instruction follow, and domain transfer. We also use a high-level search-based view (discussed in Sec. 4.2) to help systematically understand the essence of automatic jailbreak.
Furthermore, our research explored the entire pipeline of risk taxonomy, attack method, evaluation, and defense, providing a great perspective beyond prior studies that focus on only part of them. For example, Mazeika et al. (2024) emphasizes categorizing and countering harmful behaviors of LLMs through adversarial training, and Chu et al. (2024a); Zhou et al. (2024b) specifically addresses jailbreak attacks and their assessment.
Moreover, we explicitly cover multilingual (Sec. 4.1.3) and multimodal (Sec. 7) attacks against GenAI. We also discuss risks in GenAI-based applications, as generative models are integrated into workflows, such as cooperative agents or tool-enhanced agents (Naihin et al., 2023; Ruan et al., 2023) (Sec. 8).
@@ 3 Risk Taxonomy
Generative AI can exhibit harmful behaviors when manipulated by malicious queries. To assess the safety of GenAI, existing studies often analyze the models’ responses to such queries, with each query targeting a particular risk area. These studies typically focus on ethical or social risks, categorizing them based on the types of harm they could potentially cause. However, there are alternative works to categorize risks. In this chapter, we first explore how risks associated with LLMs are categorized and characterized, offering a comprehensive understanding of the landscape.
@@
To ensure that LLM-based applications are used legally and safely, users usually need to accept the usage policy that prohibits risky behaviors. We mainly focus on Meta LLM user policy and OpenAI usage policy (Meta, 2023a; OpenAI, 2024). Both organizations have the following policy in common: (1) Complying with laws. Users should not use their products to engage in illegal activities such as violence, exploitation, and terrorism. (2) Harming individuals. This includes engaging or assisting in activities that can physically harm you or others. OpenAI further requests users to not exploit their services to harm others or circumvent safeguards. Meta requests users to disclose the risks of AI systems that have used their model. Based on these policies, some work develops their own benchmarks or risk taxonomy. For example, Qi et al. (2023b) extract 11 risk categories from OpenAI and Meta’s prohibited use cases and construct harmful instructions for each category.
@@
Risks vary greatly and can lead to a diverse range of hazards, making it essential to categorize them based on the types of harm they can cause. Weidinger et al. (2021) categorize the risks associated with LLMs into six distinct areas. Building upon this, Wang et al. (2023g) propose a risk taxonomy that categorizes each risk area into more fine-grained types of harm. This taxonomy has covered most of the risks shown in other works. Mazeika et al. (2024) divide risks into 7 harm types based on their semantics (i.e., if they are semantically similar). Vidgen et al. (2023) propose a benchmark that covers 5 harm areas based on the severity and prevalence of their effects. Hua et al. (2024) define the scope of safety issues into 7 categories.
@@
LLM risks often target specific people or groups. Therefore, a taxonomy based on risk targets can be proposed. Derner et al. (2023) consider the security risks of LLMs and categorize them into user-targeted, model-targeted, and third-party-targeted. Derczynski et al. (2023) recognize 5 roles that can be at risk, including model providers, developers, text consumers, publishers, and external groups. Kirk et al. (2023) investigate the benefits and risks of personalized LLMs and divide them into individual and societal levels.
@@
Some works focus on a specific domain and propose a risk taxonomy for that domain. Tang et al. (2024) explore risks in science. They classify risks by scientific domains and impacts on outside environments. Min et al. (2023) investigate the copyright risks of LLMs and divide them by the licenses of training data: including no restrictions, permissive licenses, attribution licenses, and all others as non-permissive.
@@
Sun et al. (2023) define 8 safety scenarios when using LLMs. Each comes with a definition and its corresponding examples. Zhang et al. (2023f) drive 7 safety problems from these safety scenarios. Wang & Tong (2023) introduce risks based on scenarios of life, study, and work.
It should be noted that for different taxonomy criteria, a malicious prompt can be assigned different risks. Also, combining different risks results in different user queries. Figure 3 shows examples of prompts with multiple risks.
Figure 3: Illustration of risk taxonomy examples. We categorized and characterized the methods for assessing the risk associated with AI from five aspects. Color-coded lines connect specific examples of risky query attacks, presented below, to illustrate how each is categorized according to the relevant taxonomy criteria.
4 Attack
Attacking language models is a search problem: in the exponential-sized language space, find prompts that induce malicious behavior. The problem has been approached from two perspectives.
On the one hand, people have discovered by trial and error (Inie et al., 2023) many strategies (Sec. 4.1) of attacking language models. Examples include adding suffixes like “Sure, here is”, role play and even writing in ciphers. The strategies are diverse and often seemingly ad hoc, but we propose that these strategies are all traceable to basic capabilities and training processes of language models.
On the other hand, to find attacking prompts automatically for red teaming, searchers (Sec. 4.2) have been proposed to search the prompt space according to certain objectives. They are scalable but often lacking in diversity. In fact, these two perspectives can complement each other:search can be guided by strategies to fully explore the subspace behind the them, and new strategies can be generalized from prompts discovered by searchers.
Figure 4: Language model attack strategies by core features of language models. We classify attack strategies into four key areas: Completion Compliance, which arises from the pretraining stage and leverages the model’s text completion tendency through autoregressive generation; Instruction Indirection, utilizing capabilities enabled by instruction tuning; Generalization Glide, reflecting the model’s capacity for generalization, including reasoning and multilingualism; and Model Manipulation, which indicates a customizable model achieved by directly changing parameters, activations, or configurations.

### v2

@@ Abstract
Generative models are rapidly gaining popularity and being integrated into everyday applications, raising concerns over their safe use as various vulnerabilities are exposed. In light of this, the field of red teaming is undergoing fast-paced growth, highlighting the need for a comprehensive survey covering the entire pipeline and addressing emerging topics. Our extensive survey, which examines over 120 papers, introduces a taxonomy of fine-grained attack strategies grounded in the inherent capabilities of language models. Additionally, we have developed the “searcher” framework to unify various automatic red teaming approaches. Moreover, our survey covers novel areas including multimodal attacks and defenses, risks around LLM-based agents, overkill of harmless queries, and the balance between harmlessness and helpfulness. Warning: This paper contains examples that may be offensive, harmful, or biased. [^1] Up-to-date literature is available on: https://github.com/Libr-AI/OpenRedTeaming
@@ 1 Introduction
Generative artificial intelligence (GenAI), such as Large Language Models (LLMs) and Vision-Language Models (VLMs), have gained wide application in areas such as dialogue systems Ouyang et al. (2022), code completion Chen et al. (2021), and domain-specific usages Wu et al. (2023c). However, the generative nature and versatility of such models also introduce more vulnerabilities than conventional systems. GenAI can be induced to produce biased, harmful, or unintended outputs when presented with carefully crafted prompts, a phenomenon known as prompt attacks or LLM jailbreak Shen et al. (2023), which may facilitate the propagation of harmful information and malicious exploitation of applications utilizing GenAI.
To gain a comprehensive understanding of potential attacks on GenAI and develop robust safeguards, researchers have conducted studies on various red-teaming strategies, automated attack approaches, and defense methods. Several surveys have been composed to investigate the rapidly growing field of GenAI safety systematically. Although existing surveys provide valuable insights, we identify several areas for improvement. First, many surveys cover a limited scope of attack strategies and defenses compared to the rapidly growing body of research in this fast-evolving field. Their categorization of methods is also coarse-grained, and insensitive to the specific strategies proposed in different papers. Second, there lacks a unifying perspective that connects the full spectrum of safety taxonomies, attack methods, benchmarks, and defenses in GenAI safety. Third, emerging topics such as multimodal attacks, LLM-based application safety and overkill are often overlooked or only briefly covered, and there is a need for dedicated analysis of this rapidly evolving research area.
@@
  • We propose a novel framing of automated red-teaming methods as search problems, in line with which we decouple popular search methods into three components: the state space, search goal, and search operation, unlocking a larger space for the future design of automated red-teaming methods.
  • We pay distinct attention to emergent areas of GenAI safety, including multimodal attacks, overkill of harmless queries, as well as the safety of downstream applications powered by LLMs.
  • We propose several key future directions for advancing language model safety, identifying challenges in cybersecurity, persuasive capabilities, privacy, and domain-specific applications, and advocating for adaptive evaluation frameworks and advanced defenses to address these evolving risks.
@@
The remainder of this paper is structured as follows (also shown in Figure 2). Section 2 introduces key terminology and positions our work in the context of related surveys. Section 3 outlines a taxonomy of risks associated with LLMs. In Section 4, we dive into attack strategies and automated search methods for LLMs. Evaluation benchmarks and metrics are covered in Section 6. Section 5 presents an overview of defensive approaches. Section 7 and Section 8 discuss safety of multimodal models and LLM-based applications, respectively. Finally, we highlight promising future research directions in Section 9 and conclude in Section 10.
@@ 2 Background
As the rise of GenAI requires understanding of the associated risks posed by adversarial attacks and jailbreaks, this section introduces key terminology and differentiates this survey from existing work by highlighting its comprehensive coverage spanning risk taxonomies, attack methods, evaluation benchmarks, defenses, and emerging areas such as multimodal and LLM-based application safety.
@@ 2.1 Terminology
Many different terms have been employed in the field of AI safety. Here we try to define them and clarify their differences. Specifically, we elicit the features of attack terms in Figure 3.
Figure 3: The main differences in commonly-used attack terms.
@@
An adversarial attack is a technique used to introduce intentionally crafted perturbations or modifications to the input data to cause a model to produce incorrect or undesirable outputs Ren et al. (2020). These attacks can exploit vulnerabilities in the model’s training data, architecture, or decision boundaries, potentially leading to safety or security risks.
@@
Jailbreaking involves bypassing a GenAI’s safety measures, often through prompt engineering, to override guardrails and potentially enable harmful activities like promoting hate speech or criminal acts Chao et al. (2023); Derner & Batistic (2023). The difference between an adversarial attack and jailbreak is subtle: while the primary goal of an adversarial attack is to trigger erroneous model outputs like wrong classification results, a jailbreaker aims to obtain output that breaches established safety guidelines Shayegani et al. (2023b). Furthermore, adversarial attacks can be performed either in the training or inference phases, while jailbreaks typically happen during inference.
@@
Prompt injection, similar to SQL injection, refers to the process of inserting malicious or manipulative content into the prompts or inputs provided to a model Abdelnabi et al. (2023), and hijacking the model’s intended task. For example, the attacker may trick the model into submitting sensitive information to a malicious URL, or providing wrong summary results or false labels.
@@
Red teaming is a practice of simulating malicious scenarios to identify vulnerabilities and test the robustness of systems or models Abbass et al. (2011). Distinct from hacking or malicious attacks, red teaming in AI safety is a controlled process typically conducted by model developers that deliberately probes and manipulates models to uncover potential harmful behaviors. This process helps uncover weaknesses and informs the development of mitigation strategies. In AI safety, red teaming can employ methods such as adversarial attacks, jailbreaking, and prompt injection. Typically, major LLM companies such as OpenAI and Anthropic have their own red-teaming networks.
@@
Alignment refers to the process of ensuring that the behavior and outputs of AI are consistent with human values and ethical principles Askell et al. (2021); Ouyang et al. (2022), such as fine-tuning or RLHF discussed in Section 5.1. There are numerous alignment criteria Kiritchenko & Mohammad (2018); Carranza et al. (2023); Kirchner et al. (2022), while the common alignment goal follows the HHH criteria Askell et al. (2021):
@@
Overkill describes a language model refusing to answer a harmless question containing seemingly dangerous keywords, e.g., “How to kill a python process” or “Where can I shoot a good photo?”. This is usually a side-effect of safety alignment, as the model leans toward harmlessness instead of helpfulness in the multi-objective optimization process, compromising its practical value.
@@
Different studies have varying definitions of ASR, mainly diverging in their choice of indicator function $I(\cdot)$. For instance, Zou et al. (2023b) determines ASR based on the presence of specific keywords such as “I am sorry.” In contrast, work like Pi et al. (2024); Shu et al. (2024) employs language models to judge the success of an attack. For a detailed introduction to ASR, see Section 6.1.2.
@@ 2.2 Related Work
Our survey, taking a vast amount of prior honorable work Dong et al. (2024b); Das et al. (2024); Shayegani et al. (2023b); Deng et al. (2023c); Esmradi et al. (2023) discussing risk taxonomies into consideration, provides one of the most comprehensive and structured views of attacks and risks, based on different abilities of GenAI such as in-context learning, autoregressive modeling, instruction follow, and domain transfer. We also use a high-level search-based view (discussed in Section 4.2) to help systematically understand the essence of automatic jailbreak.
Furthermore, our research explores the entire pipeline of risk taxonomy, attack method, evaluation, and defense, providing a holistic perspective beyond prior studies that focus on only one aspect. For example, Mazeika et al. (2024) emphasizes categorizing and countering harmful behaviors of LLMs through adversarial training, and Chu et al. (2024a); Zhou et al. (2024b) specifically addresses jailbreak attacks and their assessment.
Moreover, we explicitly cover multimodal attacks against GenAI (Section 7) and discuss risks in GenAI-based applications as generative models are integrated into workflows, such as cooperative and tool-enhanced agents (Section 8) Naihin et al. (2023); Ruan et al. (2023).
@@ 3 Risk Taxonomy
Generative AI can exhibit harmful behaviors based on targeted prompting. To assess the safety of GenAI, existing studies often analyze models’ responses to such queries, with each query targeting a particular risk area. These studies typically focus on ethical or social risks, categorizing them based on the types of harm they could potentially cause. However, there have been alternative approaches to categorizing risks, as reviewed in this section.
@@
To ensure that LLM-based applications are used legally and safely, users usually need to accept an usage policy that prohibits risky behaviors. Here, we analyze the LLM usage policies of Meta and OpenAI Meta (2023a); OpenAI (2024). Both organizations share key policy guidelines: (1) Legal compliance—users must not use their products for illegal activities, including violence, exploitation, or terrorism; (2) Protection from harm—prohibiting engagement in activities that could cause physical harm to oneself or others OpenAI further requests users to not exploit their services to harm others or circumvent safeguards, while Meta requests disclosure of risks associated with AI systems using their models. Based on these policies, some work develops their own benchmarks or risk taxonomy. For example, Qi et al. (2023) extract 11 risk categories from OpenAI and Meta’s prohibited use cases and construct harmful instructions for each category, see Figure 4 (a).
Figure 4: Harm type categories in selected work.
@@
Risks vary greatly and can lead to a diverse range of hazards, making it essential to categorize them based on the types of harm they can cause. In an early study, Weidinger et al. (2021) structured the LLM risk landscape into 6 broad areas (Figure 4 (b)). More recently, Mazeika et al. (2024) created a benchmark spanning 8 categories (Figure 4 (c)), and Vidgen et al. (2023) proposed a benchmark covering five harm areas based on severity and prevalence (Figure 4 (d)). While these approaches highlight critical and focused types of harm, they are not comprehensive, and their categories often overlap. Expanding on Weidinger et al. (2021), Wang et al. (2023g) introduced a three-tier risk taxonomy with a more granular classification: 5 first-tier risk areas, 12 second-tier types, and 61 specific harm types. This taxonomy, one of the most detailed to date, covers most risks identified in other work (Figure 4 (e)).
@@
LLM risks often target specific people or groups, motivating a taxonomy based on risk targets. Derner et al. (2023) consider the security risks of LLMs and categorize them into user-targeted, model-targeted, and third-party-targeted. Derczynski et al. (2023) recognize 5 roles that can be at risk, including model providers, developers, text consumers, publishers, and external groups. Kirk et al. (2023a) investigate the benefits and risks of personalized LLMs and divide them into individual and societal levels.
@@
Some work has focused on a specific domain and proposed a risk taxonomy for that domain. Tang et al. (2024) explore risks in science. They classify risks by scientific domains and impacts on external environments. Min et al. (2023) investigate the copyright risks of LLMs and divide them by the licenses of training data: including no restrictions, permissive licenses, attribution licenses, and all others as non-permissive. Hua et al. (2024) investigate LLM agent safety issues across 6 domains, including: housekeeping, finance, medicine, food, and chemistry.
@@
Sun et al. (2023) define 8 safety scenarios when using LLMs. Each comes with a definition and corresponding examples. Zhang et al. (2023f) drive 7 safety problems from these safety scenarios. Wang & Tong (2023) introduce risks based on scenarios of life, study, and work.
It is worth noting that a malicious prompt can present different risks spanning different taxonomic criteria. Figure 5 illustrates examples of prompts that encompass multiple risks.
Figure 5: Illustration of risk taxonomy examples. We categorize the methods for assessing the risk associated with AI from five aspects. Color-coded lines connect specific examples of risky query attacks, presented below, to illustrate how each is categorized according to the relevant criteria.
Section Summary
Establishing a risk taxonomy is a foundational step in safety research, guiding most studies to focus on specific aspects of model safety. While some approaches are generalizable across various risk categories, much existing work hones in on particular risk types. In the following section, we delve into widely-used attack, defense, and evaluation methods in the literature, offering insights into practical approaches for addressing these risks.
4 Attacking Large Language Models
Attacking language models can be seen as a search problem: navigating the vast language space to find prompts that provoke abnormal behavior. The problem has been approached from two perspectives. On the one hand, through trial and error Inie et al. (2023), various strategies for attacking language models have been discovered (Section 4.1). Examples include adding suffixes like “Sure, here is”, role play, and even writing in ciphers. These strategies may appear diverse and ad hoc, yet we suggest that these strategies are all traceable to basic capabilities and training processes of language models. On the other hand, to find attacking prompts automatically for red teaming, searcher strategies have been proposed to search the prompt space according to certain objectives (Section 4.2). While scalable, these methods often lack diversity. In practical scenarios, these two approaches can be complementary: search can be guided by strategies to fully explore relevant subspaces, and new strategies can be generalized from prompts discovered by searchers.
Figure 6: Language model attack strategies based on core features of language models. We classify attack strategies into four key areas: Completion Compliance, which arises from the pretraining stage and leverages the model’s text completion tendency through autoregressive generation; Instruction Indirection, using capabilities enabled by instruction tuning; Generalization Glide, reflecting the model’s capacity for generalization, including reasoning and multilingualism; and Model Manipulation, which indicates a customizable model achieved by directly changing parameters, activations, or configurations.